# Prompting with Divide-and-Conquer Program Makes Large Language Models Discerning to Hallucination and Deception

### v4

@@ 6 Discussions and Limitations
Comparison with IO-Prompting: Superiority in Expressive Power As we proved in Sec. 4, Compared to IO-prompting, our proposed method has stronger expressive power and thus can solve harder problems.
Comparison with CoT and EoT: Disentangling the task decomposition and task resolution Compared to the prompting family of CoT and EoT, our proposed method explicitly separate the task decomposition stage and task resolution stage. Therefore, we can acquire explicit decomposed sub-task rather than intermediate thoughts proposed during decoding. Consequently, we can explicitly enumerate all sub-tasks output by the decomposition module and avoid the model from missing important sub-tasks.
Comparison with LtM and Decomposed Prompting: Parallel Sub-task Handler and Sequential Sub-task Handler Similar as our proposed method, some program-guided prompting like LtM and Decomposed Prompting also explicitly separate the task decomposition stage and task resolution stage. However, they are mainly designed for multi-step reasoning for complex tasks. Thus, they sequentially tackle the sub-tasks and assembly the resolutions. As a result, they tend to follow the flow of the deceptive contents, leading to proneness to deceptive content. More details about this comparison can be found in Appendix A.4.
Although our proposed method DaC surpasses the baselines on the proposed tasks, it still has some limitations. The first issue is that the appliance scope of DaC is still limited. More specifically, CoT, EoT, LtM and DaC are based on different algorithmic paradigms, learning to different Appliance Scopes. As pointed out by Feng et al., CoT and LtM can be considered as a neural dynamic programming algorithm. Thus, CoT is more suitable for tasks that can be bridged to dynamic programming, such as multi-step question answering. Differently, EoT is based on exploration and search, which is more suitable for planning and search, such as Game of 24 (Yao et al., 2023). Our proposed method is based on Divide-and-Conquer algorithm. Thus, it is more suitable for tasks that can be decomposed to a series sub-tasks that are disjoint or only slightly overlapped. Our future work will focus on further expand the appliance scope of DaC to more areas like question answering.
7 Conclusions
In this paper, we proposed a novel Divide-and-Conquer program-guided problem solving strategy. To guide large language models to tackle tasks requiring large quantity of repetitive sub-tasks and/or deceptive contents, DaC disentangle the processes of task decomposition, sub-task resolution and resolution assembly. In this way, we avoid the intertwined processes from interrupting each other and alleviate the intermediate errors in generating the resolution path. The experimental results show that such improvement leads to better performance on a wide range of tasks such as large integer multiplication, hallucination detection and misinformation detection. Moreover, theoretic analysis reveals that our proposed method help guide LLM to expand the expressive power to surpass original Transformers.
Appendix A Appendix
A.1 Proof to Theorem 4.2
@@ Appendix A Appendix
To construct a MLP that can simulate a fixed-size logic circuit without recurrent structure, we apply the circuit serialization in (Merrill & Sabharwal, 2023b) which order the gates based on topological order. In this way, we can represent the circuit as a sequence GATE[1], GATE[2], GATE[3],…,GATE[L], where each GATE[i]’s input only contains the output of the previous gates and the original input ${\bf x}$. Therefore, we can construct a $2L$-layer MLP base on the above serialization. Specifically, the $2i$-th and $12i+$ 1-th layers of the MLP will simulate the $GATE[i]$ as well as copy all previous inputs with activation function and concatenate them together. This can be done by concatenate an identical matrix on the GATE’s weight vector ($w_{A}$, $w_{O}$ or $w_{N}$). In this way, we can construct a MLP that precisely simulate the circuit. Since every time we concatenate the output of a gate with the input of it, the input dimension number of the final layer can be bounded by $O(|{\bf x}|+L)$. In the worst case, for a circuit of size $L$, we needs $2L$ layers to precisely simulate it. However, in many cases, a lot of gates in the circuits can be run parallelly. In such cases, the MLP could be much more shallow.
@@
After preparing the inputs, we start to construct our Transformer. The transformer first attach the position index for each token (positional embedding). After that, the inputs are forwarded into a transformer with depth of 2. Each transformer layer contains a multi-head attention layer followed by a MLP. As proved by (Merrill & Sabharwal, 2023b; Feng et al., 2023), the attention layer of Transformer can retrieve the feature of tokens whose positional embeddings satisfy specific conditions. For multi-head attention, different heads can retrieve tokens with different conditions. In the following construction, we will use this conclusion to construct attention heads with different functions.
@@
A.2 Prompting Details of DaC
@@
A.3 Decomposed Prompting and Least to Most
@@
A.4 More Discussions on Sequential Sub-task Tackling and Parallel Sub-task Tackling

### v6

@@ A.1 Discussions and Limitations
Comparison with IO-Prompting: Superiority in Expressive Power As we proved in Sec. 4, Compared to IO-prompting, DaC has stronger expressive power and thus can solve harder problems.
Comparison with CoT and EoT: Disentangling the task decomposition and task resolution Compared to the prompting family of CoT and EoT, DaC explicitly separate the task decomposition stage and task resolution stage. Therefore, we can acquire explicit decomposed sub-task rather than intermediate thoughts proposed during decoding. Consequently, we can explicitly enumerate all sub-tasks output by the decomposition module and avoid the model from missing important sub-tasks.
Comparison with LtM and Decomposed Prompting: Parallel Sub-task Handler and Sequential Sub-task Handler Similar as DaC, some program-guided prompting like LtM and Decomposed Prompting also explicitly separate the task decomposition stage and task resolution stage. However, they are mainly designed for multi-step reasoning for complex tasks. Thus, they sequentially tackle the sub-tasks and assembly the resolutions. As a result, they tend to follow the flow of the deceptive contents, leading to proneness to deceptive content.
Although DaC DaC surpasses the baselines on the proposed tasks, it still has some limitations. The first issue is that the appliance scope of DaC is still limited. More specifically, CoT, EoT, LtM and DaC are based on different algorithmic paradigms, learning to different Appliance Scopes. As pointed out by Feng et al., CoT and LtM can be considered as a neural dynamic programming algorithm. Thus, CoT is more suitable for tasks that can be bridged to dynamic programming, such as multi-step question answering. Differently, EoT is based on exploration and search, which is more suitable for planning and search, such as Game of 24 (Yao et al., 2023). DaC is based on Divide-and-Conquer algorithm. Thus, it is more suitable for tasks that can be decomposed to a series sub-tasks that are disjoint or only slightly overlapped. Our future work will focus on further expand the appliance scope of DaC to more areas like question answering.
A.2 Proof to Theorem 4.2
@@
To construct a MLP that can simulate a fixed-size logic circuit without recurrent structure, we apply the circuit serialization in (Merrill and Sabharwal, 2023a) which order the gates based on topological order. In this way, we can represent the circuit as a sequence GATE[1], GATE[2], GATE[3],…,GATE[L], where each GATE[i]’s input only contains the output of the previous gates and the original input ${\bf x}$. Therefore, we can construct a $2L$-layer MLP base on the above serialization. Specifically, the $2i$-th and $12i+$ 1-th layers of the MLP will simulate the $GATE[i]$ as well as copy all previous inputs with activation function and concatenate them together. This can be done by concatenate an identical matrix on the GATE’s weight vector ($w_{A}$, $w_{O}$ or $w_{N}$). In this way, we can construct a MLP that precisely simulate the circuit. Since every time we concatenate the output of a gate with the input of it, the input dimension number of the final layer can be bounded by $O(|{\bf x}|+L)$. In the worst case, for a circuit of size $L$, we needs $2L$ layers to precisely simulate it. However, in many cases, a lot of gates in the circuits can be run parallelly. In such cases, the MLP could be much more shallow.
@@
After preparing the inputs, we start to construct our Transformer. The transformer first attach the position index for each token (positional embedding). After that, the inputs are forwarded into a transformer with depth of 2. Each transformer layer contains a multi-head attention layer followed by a MLP. As proved by (Merrill and Sabharwal, 2023a; Feng et al., 2023), the attention layer of Transformer can retrieve the feature of tokens whose positional embeddings satisfy specific conditions. For multi-head attention, different heads can retrieve tokens with different conditions. In the following construction, we will use this conclusion to construct attention heads with different functions.
@@
A.3 Justification to Proposition 4.4
Suppose that the LLM is auto-regressively decoding $n$ tokens from an input context window with length of $C$. Then the decoding window of the $i$-th token is $C+i-1$. Thus, the average window size will be:
$\frac{\sum_{i=1}^{n}(C+i-1)}{n}=\frac{C+n-1}{2}$ (7)
Thus, when we sequentially decode all the sub-task resolutions, the total length of the decoded sequence will be $\sum_{i=1}^{k}r_{i}$. Thus the average window size will be:
$C+\frac{\sum_{i=1}^{k}r_{i}-1}{2}$ (8)
Meanwhile, when we apply Divide-and-Conquer, we parallely decode each sub-task’s resolution. Thus, for each sub-task, total window size will be $C\sum_{j=1}^{k}r_{j}+\sum_{i=1}^{k}\frac{(r_{i}-1)r_{i}}{2}$. Thus the average window size will be $C+\sum_{i=1}^{k}\frac{(r_{i}-1)r_{i}}{2\sum_{j=1}^{k}r_{j}}$. Meanwhile, with Jensen inequility, we have:
$\sum_{i=1}^{k}(r_{i}-1)r_{i}<\sum_{i=1}^{k}(r_{i}-0.5)^{2}\leq(\sum_{i=1}^{k}(r_{i}-0.5))^{2}\leq(\sum_{i=1}^{k}r_{i}-0.5k)^{2}$ (9)
Thus, when $k\geq 2$, we have:
$\sum_{i=1}^{k}(r_{i}-1)r_{i}<(\sum_{i=1}^{k}r_{i}-1)^{2}$ (10)
Thus, we have:
$C+\sum_{i=1}^{k}\frac{(r_{i}-1)^{2}}{2\sum_{j=1}^{k}r_{j}}<C+\frac{\sum_{i=1}^{k}r_{i}-1}{2}$ (11)
A.4 Prompting Details of DaC
@@
A.5 Decomposed Prompting and Least to Most
@@
A.6 Typical Tasks that Satisfy and Dissatisfy the Proposed Conditions
To better assist the prompt engineering on different tasks, we list the typical tasks that satisfy and dissatisfy the proposed conditions. In common tasks, the following tasks satisfy the proposed conditions. For such tasks, searching good decomposition prompt for DaC is likely to be helpful for the performance:
  1. 1.
    Multiplication
  2. 2.
    Fact Verification on Long Text
  3. 3.
    Auto Evaluation on Long Text
  4. 4.
    Article-level Summary
The following tasks typically do not satisfy the proposed conditions. For such tasks, searching good decomposition prompt for DaC is not very likely to be helpful for the performance:
  1. 1.
    Addition: It is too simple and violate the condition 1
  2. 2.
    Division: It does not contain parallel sub-tasks, thus violate condition 2
  3. 3.
    Multi-Round Question-Answering: It is a typical sequential task, thus violate condition 2
  4. 4.
    Planning: It is a typical sequential task, thus violate condition 2
A.7 More Discussions on Sequential Sub-task Tackling and Parallel Sub-task Tackling